OpenStudy (anonymous):

How do I find the derivative of f(x)=5x+9 at x=2?

OpenStudy (jack1):

the derivative is the same no matter what the point, and the derivative is the slope of the line, so maybe it meant "find the slope of the line at x=2"

OpenStudy (anonymous):

Really? We're studying the derivatives with limits so I thought it would have something to do with that.

OpenStudy (anonymous):

The derivative of a constant is alway zero, you can use the an^(a-1) rule to find that the derivative of that line everywhere is 5.

OpenStudy (jack1):

so the derivative of the eqn 5x + 9 is just 5 so the slope of the line at any point is 5 so at x=2, slope = 5

OpenStudy (anonymous):

Oh, great! Thank you

OpenStudy (jack1):

ooo, shiny, thanks!

OpenStudy (jack1):

you get how to find the derivative of an eqn tho...? that;s the important bit

OpenStudy (anonymous):

Yes! Just the slope, right?

OpenStudy (jack1):

yeah... but what about if f(x) = 2x^2 +10x so whats the slope at x=2...?

OpenStudy (anonymous):

That's actually similar to the next question on my assignment! Haha but that's a good question.. Is there a certain formula I can use like mitodoteira said?

OpenStudy (jack1):

yeah, for the derivative, what i do is: \[f(x) = 2x ^{2} + 10x\] so derivative is: \[f'(x) = 2\times(2x ^{(2-1)}) + 1\times(10^{1-1})\]

OpenStudy (anonymous):

Why did you use those exponents?

OpenStudy (jack1):

so take whatever x is to the power of... and times it in front then minus one from that power see how the 2 comes down and you multiply it by the other 2 and the x^2 becomes x^1 then the 1 comes down and you multiply it by the 10 and the x^1 becomes x^0... which equals 1 so it just becomes : f'(x) = 4x + 10

OpenStudy (jack1):

chose those powers just randomly

OpenStudy (jack1):

so at x= 2, slope = 4x +10 = 4*2 + 10 = 18

OpenStudy (jack1):

now you try on f(x) = 3x^6 +10x^4

OpenStudy (anonymous):

So for \[-12x^2+9x\] at x=6, it would be (i'm having difficulty figuring out the equation option) 2*(12x^(2-1))+1*9^(1-1) ?

OpenStudy (anonymous):

I just decided to use the equation that I'm trying to figure out-sorry!

OpenStudy (jack1):

yep that's it so -24x plus 9 is the slope then sub in x=6

OpenStudy (anonymous):

Then, it would be 144x+9?

OpenStudy (anonymous):

Oh, I thought I was supposed to square it.

OpenStudy (jack1):

wait... what?

OpenStudy (jack1):

no, the exponent is timeed by the number in front

OpenStudy (anonymous):

Okay! I understand now. So then it would be (-24*6)+9=-135

OpenStudy (jack1):

so if it's x^a derivative is ax^(a-1) so 2*(12x^(2-1))+1*9^(1-1)

OpenStudy (anonymous):

That makes sense now.

OpenStudy (jack1):

yep, u got it just quickly try the derivative of this: f(x) = 3x^6 +10x^4

OpenStudy (anonymous):

6(3x^(6-5)+1(9^(4-4)) Is this right so far?

OpenStudy (anonymous):

18x+9, so the derivative would be 18 if it's not at a certain place?

OpenStudy (jack1):

not quite... f(x) = 3x^6 +10x^4 so f'(x) = 6*3 x^(6-1) + 4*10 x^(4-1)

OpenStudy (jack1):

so 18x^5 +40x^3

OpenStudy (anonymous):

I thought I had to make them both even out to 1 and zero, but I see that now. So no matter what the exponents are, I always subtract 1 from them in that equation?

OpenStudy (jack1):

yep there u go

OpenStudy (anonymous):

Thank you so much! Is there anything special I have to do for fractions? For example, \[f(x)=-11/9\] at 9? Also, if there are no powers, or possibly fractions involved, the derivative is just the slope?

OpenStudy (jack1):

there's no x in that eqn...?

OpenStudy (anonymous):

-11/9x

OpenStudy (jack1):

like: \[\frac{ -11 }{ 9x } or (\frac{ -11 }{ 9 })x\]

OpenStudy (anonymous):

Sorry, I misread it. It was (-11/x) at x=9

OpenStudy (jack1):

ah, ok so 1/ anything is the same as anything to the power of -1 eg \[\frac{ 1 }{ 5 } = 5^{-1}\]

OpenStudy (jack1):

so \[\frac{ 11 }{ x } = 11 x ^{-1}\]

OpenStudy (anonymous):

Yes.. would that make the derivative 11, or do we still have to work with the exponent?

OpenStudy (jack1):

still need to work with the exponent so (-1 - 1) = ???

OpenStudy (anonymous):

-2

OpenStudy (jack1):

so it becomes (-1) * (-11) *(x^-2)

OpenStudy (jack1):

follow ok...?

OpenStudy (jack1):

more explainin...?

OpenStudy (anonymous):

Yeah, I'm just not sure what to do after that other than solve, and in decimal form, I got -0.00826, but I feel like that is incorrect.

OpenStudy (jack1):

9^2 = 81 so 11/81 = 0.1358

OpenStudy (anonymous):

Wow, that was a silly mistake of mine. I got it now! Thank you so much

OpenStudy (jack1):

always welcomes